# MultiBanana: A Challenging Benchmark for Multi-Reference Text-to-Image Generation

### v1

@@ Abstract
Recent text-to-image generation models have acquired the ability of multi-reference generation and editing; the ability to inherit the appearance of subjects from multiple reference images and re-render them under new contexts. However, the existing benchmark datasets often focus on the generation with single or a few reference images, which prevents us from measuring the progress on how model performance advances or pointing out their weaknesses, under different multi-reference conditions. In addition, their task definitions are still vague, typically limited to axes such as “what to edit” or “how many references are given”, and therefore fail to capture the intrinsic difficulty of multi-reference settings. To address this gap, we introduce MultiBanana, which is carefully designed to assesses the edge of model capabilities by widely covering multi-reference-specific problems at scale: (1) varying the number of references, (2) domain mismatch among references (e.g., photo vs. anime), (3) scale mismatch between reference and target scenes, (4) references containing rare concepts (e.g., a red banana), and (5) multilingual textual references for rendering. Our analysis among a variety of text-to-image models reveals their superior performances, typical failure modes, and areas for improvement. MultiBanana will be released as an open benchmark to push the boundaries and establish a standardized basis for fair comparison in multi-reference image generation. Our data and code are available at https://github.com/matsuolab/multibanana.
Figure 1: The overview of MultiBanana. MultiBanana widely covers multi-reference specific problems, varying the number of references (the top row), domain and scale mismatch among references (two on the left in the middle row), multilingual text rendering (center in the bottom row), and containing rare concepts (right in the bottom row).
@@ 1 Introduction
Recent advances in image generation, driven by large-scale multimodal LLM backbones, have enabled unprecedented levels of instruction-following [google2025nanobanana, openai2025gpt4oimage, wu2025qwenimagetechnicalreport, cao2025hunyuanimage3, wu2025less, xiao2024omnigen, wu2025omnigen2, xia2025dreamomni, xia2025dreamomni2, cui2025emu35nativemultimodalmodels]. Among these developments, multi-reference image generation has emerged as a particularly notable capability. When a user provides multiple reference images, the model can inherit the subject’s appearance and re-render it in a new context while faithfully preserving identity. This shift moves image generation beyond single-reference conditioning toward greater controllability, driving increasing demand in industrial domains such as content production [ruiz2022dreambooth, ye2023ip-adapter], advertising [horita2024retrievalaugmented, morita2025tkg], and fashion design [zhu2023tryondiffusion, choi2024improving, fang2024vivid], where reference-guided generation is essential.
However, despite this progress, we lack a benchmark that systematically evaluates how model performance varies under different multi-reference conditions. As shown in Table 1, existing benchmarks [wang2023editbench, basu2023editval, sheynin2024emuedit, zhang2023magicbrush, huang2024smartedit, yu2025anyedit, ma2024i2ebench, ye2025imgedit, ruiz2022dreambooth, wu2025omnigen2, xia2025dreamomni2] suffer from two fundamental limitations. First, they have an extremely narrow coverage due to the small number of unique reference images, making it impossible to assess the impact of varying reference conditions in both single- and multi-reference tasks. Second, their task definitions still lack details, typically limited to axes such as“what to edit” or “how many references are given”, and therefore fail to capture the intrinsic difficulty of multi-reference settings. Moreover, existing benchmarks limit the number of references to a narrow range (typically 1–5) and do not examine how performance changes with more references.
To address these limitations, we introduce a new evaluation framework grounded not only in task categories but also in the intrinsic properties of the reference set itself. Our benchmark incorporates multiple orthogonal axes specific to the multi-reference regime, including broader coverage of the number of references (up to 8), domain gaps between references (e.g., photo–anime mixtures), mismatches in object scale between reference and target scenes, rare concepts in the references (e.g., a red banana), and multilingual textual rendering references (English, Chinese, and Japanese).
Using this benchmark, we conduct a cross-conditional analysis to assess model capabilities, identify characteristic failure modes, and point out remaining gaps. For example, while performance degradation due to increasing reference counts can be mitigated by agent-style multi-stage generation pipelines, failures arising from unsupported styles or languages remain difficult to remedy with inference-time strategies. We will open-source MultiBanana to establish a standard foundation for fair comparison and to accelerate progress in multi-reference image generation.
@@
EditBench [wang2023editbench] 240 1 1 CLIP [radford2021clip]
EditVal [basu2023editval] 648 13 1 CLIP, VLM, manual
EmuEdit [sheynin2024emuedit] 3055 7 1 L1, CLIP, DINO [caron2021dino]
MagicBrush [zhang2023magicbrush] 1053 9 1 L1, L2, CLIP, DINO
AnyEdit [yu2025anyedit] 1250 25 1 L1, CLIP, DINO
I2EBench [ma2024i2ebench] 2240 16 1 GPT [openai2023gpt4]
ImgEdit-Bench [ye2025imgedit] 811 14 1 GPT (3 dim.), Fake Det. [xu2024fakeshield]
DreamBooth [ruiz2022dreambooth] 75 1 1 CLIP. DINO
OmniContext [wu2025omnigen2] 400 8 3 GPT (3 dim.)
DreamOmni2 [xia2025dreamomni2] 319 29 4 Gemini [geminiteam2023gemini], Doubao [bytedance2025doubao]
@@ 2.1 Controllable Text-to-Image Generation
Diffusion models have achieved state-of-the-art performance in high-fidelity image synthesis. Recent large pretrained diffusion models with textual conditioning, such as Stable Diffusion [rombach2022ldm, podell2024sdxl, esser2024sd3] and FLUX [flux2024, labs2025flux1kontext], enable the synthesis of novel scene images while preserving the identity depicted in the reference images. Building on these developments, approaches based on fine-tuning (e.g., DreamBooth [ruiz2022dreambooth]) and adapter-based methods (e.g., IP-Adapter [ye2023ip-adapter]) have been actively explored to address the controllability [jiang2025infiniteyou, gal2023an, brooks2023instructpix2pix, mokady2023null, tan2025ominicontrol, NEURIPS2023_602e1a5d]. Additionally, training-free approaches (e.g., prompt-to-prompt [hertz2023prompttoprompt]) have also been widely proposed [parmar2023zero, Cao_2023_ICCV, Huberman-Spiegelglas_2024_CVPR, Wallace_2023_CVPR, Tumanyan_2023_CVPR, kulikov2025flowedit, miyake2025negative-prompt-inversion]. Recent research on unified image generation has gained momentum, aiming to handle multiple image generation tasks within a single model. For example, ChatGPT-Image-1 [openai2025gpt4oimage] and Nano Banana [google2025nanobanana] can jointly process text and image inputs, enabling integrated image generation and editing within a unified framework. Similarly, the open-source models, such as Qwen-Image [wu2025qwenimagetechnicalreport] and FLUX.1-Kontext-dev [labs2025flux1kontext], built upon a robust multimodal large language model backbone, demonstrates high-quality and flexible image generation and editing performance. Beyond these, numerous open-source unified generative models continue to emerge [xiao2024omnigen, xia2025dreamomni, wu2025omnigen2, xia2025dreamomni2]. These advancements collectively demonstrate that diffusion models are becoming increasingly flexible and adaptable across diverse conditional generation settings.
@@ 2.2 Benchmarks for Reference-Based Generation
Among reference-driven image generation tasks, image editing can be the most widely recognized, which has referred to the task of generating an edited image based on a single reference image and a textual instruction describing how the edit should be performed. Representative benchmarks such as MagicBrush [zhang2023magicbrush] and EMU-Edit [sheynin2024emuedit] have extended task-specific evaluations but still rely heavily on similarity-based metrics. SmartEdit [huang2024smartedit] supports editing of complex scenes, but does not sufficiently cover more general settings. Similarly, I2E-Bench [ma2024i2ebench] employs GPT-4o to provide human-aligned evaluations across a variety of editing tasks, yet it uses different metrics for each task and therefore fails to capture the shared characteristics of editing as a whole. More recently, ImgEdit [ye2025imgedit] has enabled the evaluation of multi-turn editing tasks, where image generation and editing are alternated interactively. One of the earliest benchmarks for reference-based image generation beyond instruction-based editing is the DreamBooth [ruiz2022dreambooth] dataset. However, such benchmarks are still limited to generation tasks conditioned on a single reference image and thus cannot handle more complex or compositional settings. Recently, following advances demonstrated by GPT-Image-1 [openai2025gpt4oimage] and Nano Banana [google2025nanobanana], increasing attention has been directed toward multi-reference image generation, where multiple reference images are used jointly for generation. Benchmarks such as OmniContext [wu2025omnigen2] and DreamOmni2 [xia2025dreamomni2] have been proposed to evaluate this task. Nevertheless, these benchmarks encompass only a small number of images and task types, thereby limiting the scope of the evaluation. Moreover, despite being designed for multi-reference generation, they do not evaluate qualitative differences or relationships between reference images, and their setups remain similar to those of conventional single-image editing benchmarks.
@@ 3.1 Task Types
In reference-driven image generation, tasks can be categorized based on the number and relational structure of reference images. In this work, we follow the overall design philosophy of prior benchmarks such as DreamOmni2 [xia2025dreamomni2], but place a stronger emphasis on an aspect that has been largely underexplored so far: the diversity and compositional relationships among multiple references. While existing benchmarks mainly evaluate performance in terms of the number of references or single-image editing accuracy, our benchmark explicitly aims to measure a model’s ability to understand and reason about inter-reference variations as a key part of multimodal generation.
Single Reference.   This is the most fundamental setting, corresponding to the standard image editing task. Given a single reference image and a textual instruction, the model generates an edited image that maintains semantic consistency and preserves the subject’s visual identity. This configuration follows traditional instruction-driven editing benchmarks such as MagicBrush [zhang2023magicbrush] and EMU-Edit [sheynin2024emuedit], which represent the foundation of reference-driven generation.
@@
Multi References – Compositional Generation and Inter-Reference Relationships.   The multi-reference setting uses 3–8 reference images, requiring the model to perform more compositional and reasoning-based generation. While prior works, such as DreamOmni2 [xia2025dreamomni2], mainly focus on compositional generation using multiple references, our benchmark takes a step further by making inter-reference diversity, consistency, and structural relations central to the evaluation.
We define four base task structures, each evaluated with $\{3,4,5,6,7,8\}$ reference images, and include variants with and without textual instructions — resulting in a total of 48 tasks:
@@
This design enables us to assess whether a model can effectively integrate multiple reference sources into a coherent, contextually consistent scene.
Diversity Factors in Multi-Reference Generation.   A central contribution of our benchmark is the explicit incorporation of inter-reference diversity factors, which have been largely overlooked in previous benchmarks such as DreamOmni2. We highlight the following new evaluation dimensions:
  • Number of references: systematic comparison across 3–8 inputs.
@@
These factors go beyond conventional image-editing evaluation and are critical for assessing a model’s compositional reasoning, identity preservation, and cross-domain generalization. In summary, our benchmark diverges from the conventional notion of multi-reference generation, such as in DreamOmni2, which primarily focuses on simple reference composition. Instead, it introduces a new framework designed to evaluate model capabilities to interpret relationships among heterogeneous references and integrate them into coherent generations.
Figure 3: (Left) Comparison between the statistics of real data only and those after adding synthetic data. The original dataset was biased toward background images, with few person- and object-related samples. To correct this imbalance, we generated additional synthetic images using Nanobanana and ChatGPT-Image-1, focusing on clear subjects such as people, animals, and objects. This significantly increased person- and object-related categories, resulting in a more balanced and comprehensive benchmark. (Right) Examples of synthesized images in each category.
@@ 3.2 How to Construct MultiBanana
Image Collection.   The reference images used in our benchmark consist of both real and synthetic data. All real images were extracted from the LAION-5B dataset [schuhmann2022laion]. To ensure high quality, we selected only images with an Aesthetic score higher than 6.25 and a resolution greater than 512 pixels. However, the collected real images exhibited a distributional bias (Figure 3; left). Specifically, landscape scenes appeared frequently, whereas categories useful as editing references, such as humans and objects, were underrepresented. While landscape images can serve as useful background references, the scarcity of human and object references limits the diversity of editing tasks that can be constructed in the benchmark. To address this imbalance, we generated additional synthetic images under diverse conditions using Nano Banana [google2025nanobanana] and ChatGPT-Image-1 [openai2025gpt4oimage], and used them as supplementary reference images (Figure 3; right). This approach not only mitigated the categorical bias in the reference set but also expanded the coverage and diversity of reference images by incorporating both real and synthetic examples.
@@
Image Filtering.  Among the collected images, some contained small subjects that were difficult to use as references, or were inappropriate as reference images, such as those containing harmful content, tables, or corrupted data. To filter out such images and retain only those suitable for a reference-based image editing benchmark, we applied the following procedure. Following [ye2025imgedit], we extracted images that contained large and clearly visible objects suitable for editing tasks. Object detection was performed using YOLOv12 [tian2025yolov12] and SAM2 [ravi2025sam], while CLIP [radford2021clip] was used to verify semantic consistency. Additionally, harmful or corrupted images were removed through a combination of automatic filtering with Gemini and manual inspection, resulting in the final curated set of real images.
Figure 4: (Left) Breakdown of the multi-reference tasks. The editing sets were selected to ensure that the number of sets within each task is balanced across different reference counts. (Middle) For every X–references task, the dataset contains at least 390 editing sets. Further, each colored task category also includes at least 70 sets, which is larger than the prior work [xia2025dreamomni2]. (Right) Word cloud generated from all prompts. It primarily consists of terms that describe a wide range of object categories as well as words indicating spatial directions.
@@
Category Classification. To construct multi-reference image generation tasks, we categorized each reference image accurately. For example, in a portrait transformation task that modifies a person’s hairstyle or makeup, both reference images must contain a person. Similarly, in replacement tasks involving humans or objects, both reference images must include either a person or an object, and for background replacement tasks, images categorized as background must be used. To satisfy such task-dependent conditions, we performed hierarchical image classification on a combined set of real and synthetic images. We first obtained automatic multi-level labels using Gemini and then conducted human verification to ensure correct categorization. At the top level, we defined five major categories: person, object, background, light, tone. Each of these top-level categories was further divided into more specific subcategories. For instance, under the person category, we included male, female, hairstyle, makeup, pose useful, and face expression; and under the object category, we defined animal, clothing design, industrial design, and object. This hierarchical labeling enabled precise task construction and increased the benchmark’s diversity.
@@
Task Construction. The editing instructions were generated using Gemini. For each task, we manually selected suitable reference categories, randomly sampled images, and presented them to Gemini to produce corresponding editing instructions. Gemini then evaluated whether following each instruction would lead to visual breakdowns. For instance, in a replacement task, whether the replaced object would appear unnaturally suspended in mid-air.
Among the generated candidates, we retained only the reference–instruction pairs that did not exhibit visual breakdowns and then asked Gemini to further assess whether each instruction constituted a challenging editing scenario, such as cross-domain modifications. Finally, we manually reviewed the remaining data to remove any residual cases that Gemini failed to detect or instructions that were deemed too trivial.
@@
Difficult Reference Categorization. For the cross-domain category, we selected samples whose reference images spanned at least two domains (e.g., color photos, anime, paintings). The different-scale category included samples with both close-up and non–close-up views. The rare-concept category contained samples featuring animals with unusual colors or patterns, humans or animals with atypical makeup or clothing, or fictional creatures. The multilingual category included samples with text in two different languages among English, Chinese, and Japanese. Samples not matching any category were grouped as others.
Table 2: Difficult reference combination statistics and ratios relative to the total 3769 tasks. Our benchmark provides a sufficient number of samples for assessing a model’s compositional reasoning, identity preservation, and cross-domain generalization, compared to prior work [xia2025dreamomni2].

### v2

@@ Abstract
Recent text-to-image generation models have acquired the ability of multi-reference generation and editing; that is, to inherit the appearance of subjects from multiple reference images and re-render them in new contexts. However, existing benchmark datasets often focus on generation using a single or a few reference images, which prevents us from measuring progress in model performance or identifying weaknesses when following instructions with a larger number of references. In addition, their task definitions are still vague, limited to axes such as “what to edit” or “how many references are given”, and therefore fail to capture the challenges inherent in combining heterogeneous references. To address this gap, we introduce MultiBanana, which is designed to assess the edge of model capabilities by widely covering problems specific to multi-reference settings: (1) varying the number of references (up to 8), (2) domain mismatch among references (e.g., photo vs. anime), (3) scale mismatch between reference and target scenes, (4) references containing rare concepts (e.g., a red banana), and (5) multilingual textual references for rendering. Our analysis among a variety of text-to-image models reveals their respective performances, typical failure modes, and areas for improvement. MultiBanana is released as an open benchmark to push the boundaries and establish a standardized basis for fair comparison in multi-reference image generation. Our data and code are available at https://github.com/matsuolab/multibanana.
Figure 1: The overview of MultiBanana. MultiBanana broadly covers problems specific to multi-reference settings, including varying the number of references (top row), domain and scale mismatches among references (two on the left in the middle row), multilingual text rendering (center in the bottom row), and the presence of rare concepts (right in the bottom row).
@@ 1 Introduction
Recent advances in image generation, driven by large-scale multimodal LLM backbones, have enabled unprecedented levels of instruction-following [22, 56, 79, 7, 81, 85, 80, 84, 83, 13]. Among these developments, multi-reference image generation has emerged as a notable capability. When a user provides multiple reference images, the model can inherit the subject’s appearance and re-render it in a new context while faithfully preserving identity. This shift moves image generation beyond single-reference conditioning toward greater controllability, driving increasing demand in industrial domains such as content production [66, 90, 78], advertising [33, 28, 51], and fashion design [98, 11, 16], where reference-guided generation is essential.
Despite this progress, we lack a benchmark to evaluate how well models follow instructions that involve multiple references under varying conditions. As shown in Table 1, existing benchmarks [77, 3, 69, 93, 30, 92, 47, 91, 66, 80, 83] suffer from two fundamental limitations. First, existing benchmarks limit the number of references to a narrow range (typically 1–4) and do not examine how well models perform when following instructions with a larger number of references. Second, their task definitions are still vague, limited to axes such as “what to edit” or “how many references are given”, and therefore fail to capture the challenges inherent in combining heterogeneous references.
To address these limitations, we introduce a new evaluation framework that is grounded in the core challenges of multi-reference generation. Our benchmark incorporates multiple orthogonal axes specific to the multi-reference regime, including broader coverage of the number of references (up to 8), domain gaps between references (e.g., photo–anime mixtures), mismatches in object scale between reference and target scenes, rare concepts in the references (e.g., a red banana), and multilingual textual rendering references (English, Chinese, and Japanese).
Using this benchmark, we analyze model capabilities, identify characteristic failure modes, and highlight remaining gaps. For example, our analysis reveals that as instruction difficulty increases (through more references, cross-domain mixtures, or rare concepts), closed-source models (e.g., Nano Banana) tend to incorporate all required subjects but suffer from compositional distortion due to overfitting to reference details, while open-source models (e.g., Qwen-Image) generate visually cleaner outputs but often omit some reference subjects altogether. We release MultiBanana as an open-source benchmark to establish a standard foundation for fair comparison and to accelerate progress in multi-reference image generation.
@@
EditBench [77] 240 1 1 CLIP [62]
EditVal [3] 648 13 1 CLIP, VLM, manual
EmuEdit [69] 3055 7 1 L1, CLIP, DINO [8]
MagicBrush [93] 1053 9 1 L1, L2, CLIP, DINO
AnyEdit [92] 1250 25 1 L1, CLIP, DINO
I2EBench [47] 2240 16 1 GPT [54]
ImgEdit-Bench [91] 811 14 1 GPT (3 dim.), Fake Det. [87]
DreamBooth [66] 75 1 1 CLIP. DINO
OmniContext [80] 400 8 3 GPT (3 dim.)
DreamOmni2 [83] 319 29 4 Gemini [19], Doubao [5]
@@ 2.1 Controllable Text-to-Image Generation
Diffusion models have achieved state-of-the-art performance in high-fidelity image synthesis [70, 26]. Large pretrained diffusion models with textual conditioning, such as Stable Diffusion [65, 60, 15] and FLUX [42, 41], can generate novel images while preserving the identity depicted in the reference images. Building on these developments, approaches based on fine-tuning (e.g., DreamBooth [66]) and adapter-based methods (e.g., IP-Adapter [90]) have been explored to improve the controllability [34, 18, 4, 50, 72, 44]. Recent research on multimodal LLM-based image generation has gained momentum, aiming to handle multiple image generation tasks within a single model. For example, GPT-Image-1 [56] and Nano Banana [22] can jointly process text and image inputs, enabling integrated image generation and editing within a unified framework. Similarly, open-source models such as Qwen-Image [79] and FLUX.1-Kontext-dev [41] demonstrate high-quality, flexible image generation and editing. Beyond these, numerous open-source unified generative models continue to emerge [85, 84, 80, 83]. These advancements demonstrate that diffusion models are becoming flexible and adaptable across diverse conditional generation settings.
@@ 2.2 Benchmarks for Reference-Based Generation
Among reference-driven image generation tasks, image editing is the most widely recognized: it involves generating an edited image from a single reference image and a textual instruction describing how the edit should be performed. Representative benchmarks such as MagicBrush [93] and EMU-Edit [69] have extended task-specific evaluations but still rely heavily on similarity-based metrics. SmartEdit [30] supports editing of complex scenes, but does not sufficiently cover more general settings. Similarly, I2E-Bench [47] employs GPT-4o to provide human-aligned evaluations across a variety of editing tasks, yet it uses different metrics for each task and therefore fails to capture the shared characteristics of editing as a whole. ImgEdit [91] has enabled evaluation of multi-turn editing tasks, in which image generation and editing are alternated interactively. One of the earliest benchmarks for reference-based image generation beyond instruction-based editing is the DreamBooth [66] dataset. However, such benchmarks are still limited to generation tasks conditioned on a single reference image and thus cannot handle more complex or compositional settings. Recently, advances demonstrated by GPT-Image-1 [56] and Nano Banana [22] have drawn increasing attention to multi-reference image generation, where multiple reference images are jointly used for generation. Benchmarks such as OmniContext [80] and DreamOmni2 [83] have been proposed to evaluate this task. Nevertheless, these benchmarks encompass only a small number of images and task types, thereby limiting the scope of the evaluation. Moreover, although designed for multi-reference generation, they do not account for the heterogeneous nature of reference images. This challenge is analogous to instruction following in LLMs [97, 35, 24, 23, 40], where performance degrades as the number of constraints grows; multi-reference image generation similarly requires models to satisfy all given references without neglecting any.
@@ 3.1 Task Types
In reference-driven image generation, tasks can be categorized based on the number and relational structure of reference images. In this work, we follow the overall design philosophy of prior benchmarks such as DreamOmni2 [83], but extend the number of references up to 8 to enable a more fine-grained analysis of scalability. We also go beyond task-category-based evaluation by incorporating axes that capture the intrinsic properties of the reference set itself, such as domain gaps between references, scale mismatches, rare concepts, and multilingual text rendering.
Single Reference.   This is the simplest setting, corresponding to the standard image editing task. Given a single reference image and a textual instruction, the model generates an edited image that maintains semantic consistency and preserves the subject’s visual identity. This configuration follows traditional instruction-driven editing benchmarks such as MagicBrush [93] and EMU-Edit [69], which represent the foundation of reference-driven generation.
@@
Multi References – Compositional Generation and Inter-Reference Relationships.   The multi-reference setting uses 3–8 reference images, requiring the model to perform more compositional generation. We define four base task structures, each evaluated with $\{3,4,5,6,7,8\}$ reference images, resulting in 24 tasks:
@@
This design enables us to assess whether a model can integrate multiple reference sources into a consistent scene.
Diversity Factors in Multi-Reference Generation.   A central contribution of our benchmark is the explicit incorporation of inter-reference diversity factors, which have been largely overlooked in previous benchmarks such as DreamOmni2 [83]. We highlight the following dimensions:
@@
These factors go beyond prior benchmarks that focus on simple reference composition; our benchmark evaluates whether models can interpret relationships among heterogeneous references and integrate them into coherent outputs.
Figure 3: (Left) Comparison between the statistics of real data only and those after adding synthetic data. The original dataset was biased toward background images, with few person- and object-related samples. To correct this imbalance, we generated additional synthetic images using Nano Banana and GPT-Image-1, focusing on clear subjects such as people, animals, and objects. This results in a more balanced and comprehensive distribution of data. (Right) Examples of synthesized images in each category.
@@ 3.2 How to Construct MultiBanana
Image Collection.   The reference images used in our benchmark consist of both real and synthetic data. All real images were extracted from the LAION-5B dataset [68]. To ensure high quality, we selected only images with an Aesthetic score [43] higher than 6.25 and a resolution greater than 512 pixels. However, the collected real images exhibited a distributional bias (Figure 3; left). Specifically, landscape scenes appeared frequently, whereas humans and objects were underrepresented. While landscape images can serve as useful background references, the scarcity of human and object references limits the diversity of editing tasks. To address this imbalance, we generated additional synthetic images under diverse conditions using Nano Banana [22] and GPT-Image-1 [56], and used them as supplementary reference images (Figure 3; right). This approach not only mitigated the categorical bias in the reference set but also expanded its coverage and diversity by incorporating both real and synthetic examples.
@@
Image Filtering.  Among the collected images, some contained small subjects that were difficult to use as references, or were inappropriate for reference use, such as images containing harmful content, tables, or corrupted data. To filter out such images and retain only those suitable for a reference-based image editing benchmark, we applied the following procedure. Following [91], we extracted images that contained large and clearly visible objects suitable for editing tasks. Object detection was performed using YOLOv12 [73] and SAM2 [64], while CLIP [62] was used to verify semantic consistency. Additionally, harmful or corrupted images were removed through a combination of automatic filtering with Gemini and manual inspection, resulting in the final curated set of real images.
@@
Category Classification. To construct multi-reference image generation tasks, we categorized each reference image. For example, in a portrait transformation task that modifies a person’s hairstyle or makeup, both reference images must contain a person. Similarly, in replacement tasks involving humans or objects, both reference images must include either a person or an object, and for background replacement tasks, images categorized as background must be used. To satisfy such task-dependent conditions, we performed hierarchical image classification on a combined set of real and synthetic images. We first obtained automated multi-level labels using Gemini, then conducted human verification to ensure correct categorization. At the top level, we defined six major categories: person, object, background, light, tone, and style. Each of these top-level categories was further divided into more specific subcategories. For instance, under the person category, we included male, female, hairstyle, makeup, pose useful, and face expression; and under the object category, we defined animal, clothing design, industrial design, and object. This hierarchical labeling enabled systematic task construction.
@@
Task Construction. The editing instructions were generated using Gemini. For each task, we manually selected suitable reference categories, randomly sampled images, and presented them to Gemini to generate corresponding editing instructions. Gemini then evaluated whether following each instruction would lead to visual breakdowns. For instance, in a replacement task, whether the replaced object would appear unnaturally suspended in mid-air. Among the generated candidates, we retained only the reference–instruction pairs that did not exhibit visual breakdowns and then asked Gemini and human annotators to further filter out tasks that were nonsensical or too trivial.
@@
Difficult Reference Categorization. For the cross-domain category, we selected samples whose reference images spanned at least two domains (e.g., color photos, anime, paintings). The different-scale category included samples with both close-up and non–close-up views. The rare-concept category contained samples featuring animals with unusual colors or patterns, humans or animals with atypical makeup or clothing, or fictional creatures. The multilingual category included samples with text in two different languages, including English, Chinese, and Japanese.
Figure 4: (Left) Breakdown of the multi-reference tasks. The editing sets were selected to ensure that the number of sets per task is balanced across reference counts. (Middle) For every X–references task, the dataset contains at least 390 editing sets. Furthermore, each colored task category includes at least 70 sets, which is larger than in prior work [83]. (Right) Word cloud generated from all prompts. It primarily consists of terms that describe a wide range of object categories and words indicating spatial directions.
3.3 Statistics in MultiBanana
In the real dataset, the largest major category is background, comprising 798 images, one-fourth of the total. These mainly consist of images without distinct subjects. The next-largest categories are person (741 images) and style (579 images), and together these three categories account for 70% of the dataset. In contrast, light (195 images), object (413 images), and tone (118 images) are small in number. Although these categories are useful as references for changing global conditions such as illumination and overall atmosphere, their sample sizes remain limited. Within the object category, there are 111 images of animals, 40 of clothing and textile design, 71 of industrial design, and 191 of miscellaneous objects, indicating that data focusing on specific materials or design patterns is scarce. Most images in the background, style, and tone categories lack clear subjects. Furthermore, images containing textual elements, such as logos or signs, are nearly absent. The person category is subdivided into finer attributes: face expression (10), female (181), male (139), hairstyle (135), makeup (59), and pose usefulness (217). While pose-related data account for the largest portion (around 30%), reference data capturing subtle characteristics such as facial expression or makeup remain limited, resulting in a skewed distribution.
To address this imbalance, synthetic data were generated, focusing primarily on images containing explicit subjects (e.g., persons, animals, and objects) and on those including text. The generation process was designed to intentionally expand the coverage of underrepresented subcategories and improve the overall balance among attributes.
[…]
Task Statistics.   After filtering the editing instructions using both Gemini and manual checks, we obtained 3,769 high-quality reference images and instruction sets. Among them, 264 sets correspond to single-reference tasks, 907 to two-reference tasks, and 2,598 to multi-reference tasks. Figure 4 provides a further breakdown of the multi-reference tasks, showing that they are evenly distributed with respect to the number of reference images. Additionally, for each task type, X Object, X–1 Object + Background, X–1 Object + Local, and X–1 Object + Global, we ensured that at least 70 sets were included. This is larger than the number of sets in multi-reference tasks of OmniContext [80] and DreamOmni2 [83], enabling more reliable benchmarking. Furthermore, Table 2 shows that our newly introduced evaluation axes – cross-domain, scale and viewpoint differences, rare concept, and multilingual – provide sufficient samples. See Appendix C for details.
Table 2: Statistics and ratios of difficult reference combinations relative to the total 3,769 tasks. Our benchmark provides sufficient samples for assessing a model’s capacity to interpret relationships among heterogeneous references and integrate them into outputs.